Use Simpson's Rule to estimate  ∫_0^2▒1/8 e^(x^2 ) dx with a maximum error of 0.1


1
Expert's answer
2021-05-13T12:34:49-0400

"Given, f(x)=\frac{e^{x^2}}{8}.\newline
Exact\space value, \int_0^2\frac{e^{x^2}}{8}dx=2.0\space (2 \text{one decimal place})\newline
\text{By using simpson's rule,
the approximate value is} S_4\newline
S_4=\frac{\Delta x}{3}[f(x_0)+4f(x_1)+2f(x_2)+4f(x_3)+f(x_4)]\newline
where\space x_0=0, x_1=0.5, x_2=1,x_3=1.5, x_4=2, n=4, \Delta x=\frac{b-a}{n}=\frac{2-0}{4}=0.5\newline
S_4=\frac{1}{3}[\frac{1}{8}+4(\frac{e^{0.5^2}
}{8})+2(\frac{e}{8})+4(\frac{e^{1.5^2}}{8})+
(\frac{e^4}{8})]\newline
=2.1 \space (2 \text{one decimal place})
\newline
Error\leq 0.1\newline
|exact-approx|=|2.0566-2.1692|\leq 0.1"